At 1 bar, how much energy is required to heat 79.0 g of H,O(s) at – 24.0°C to H,O(g) at 123.0 °C? q= k.
Answer #1

Ti = -24.0 oC
Tf = 123.0 oC
here
Cs = 2.09 J/goC

Heat required to convert solid from -24.0 oC to 0.0 oC
Q1 = m*Cs*(Tf-Ti)
= 79 g * 2.09 J/goC *(0--24) oC
= 3962.64 J

Lf = 333.6 J/g

Heat required to convert solid to liquid at 0.0 oC
Q2 = m*Lf
= 79.0g *333.6 J/g
= 26354.4 J

Cl = 4.184 J/goC

Heat required to convert liquid from 0.0 oC to 100.0 oC
Q3 = m*Cl*(Tf-Ti)
= 79 g * 4.184 J/goC *(100-0) oC
= 33053.6 J

Lv = 2260.0 J/g

Heat required to convert liquid to gas at 100.0 oC
Q4 = m*Lv
= 79.0g *2260.0 J/g
= 178540 J

Cg = 2.01 J/goC

Heat required to convert vapour from 100.0 oC to 123.0 oC
Q5 = m*Cg*(Tf-Ti)
= 79 g * 2.01 J/goC *(123-100) oC
= 3652.17 J

Total heat required = Q1 + Q2 + Q3 + Q4 + Q5
= 3962.64 J + 26354.4 J + 33053.6 J + 178540 J + 3652.17 J
= 245563 J
= 245.6 KJ
Answer: 245.6 KJ